OpenStudy (anonymous):

pleaseee solve 2(x – 3) – 17 = 13 – 3(x + 2)

OpenStudy (phi):

do you know how to get rid of the parentheses by multiplying the 2 times what is inside? Same for the -3 on the other side. Can you do that?

jhonyy9 (jhonyy9):

2x-6-17=13-3x-6 5x=30 x=30/5 =6 x=6